would it take 16 Venus, the second closest planet to the sun, has a surface temperature of 7.3 x 102 K. Convert this tempera- ture to degrees Fahrenheit and degrees Celsius.

please do all steps
Answer #1

the definition implies that absolute zero is equal to -273.15 °C

7.3*10^2 k = 730 k

Convert kelvin to  °C

T( °C) = T(K) - 273.15

so -273.15 + 730 = 456.85 °C

convert Kelvin to fahrenheit

T( °F) = T(K) *9/5 - 459.67

convert 730 k to f

T( °F) = 730* 9 / 5 -459.67

T( °F) = 854.33  °F